### v3

# Subgraph-Aware Training of Text-based Methods for Knowledge Graph Completion

Abstract
Fine-tuning pre-trained language models (PLMs) has recently shown a potential to improve knowledge graph completion (KGC). However, most PLM-based methods encode only textual information, neglecting various topological structures of knowledge graphs (KGs). In this paper, we empirically validate the significant relations between the structural properties of KGs and the performance of the PLM-based methods. To leverage the structural knowledge, we propose a Subgraph-Aware Training framework for KGC (SATKGC) that combines (i) subgraph-aware mini-batching to encourage hard negative sampling, and (ii) a new contrastive learning method to focus more on harder entities and harder negative triples in terms of the structural properties. To the best of our knowledge, this is the first study to comprehensively incorporate the structural inductive bias of the subgraphs into fine-tuning PLMs. Extensive experiments on four KGC benchmarks demonstrate the superiority of SATKGC. Our code is available.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
Factual sentences, e.g., Leonardo da Vinci painted Mona Lisa, can be represented as entities, and relations between the entities. Knowledge graphs treat the entities (e.g., Leonardo da Vinci and Mona Lisa) as nodes, and the relations (e.g., painted) as edges. Each edge and its endpoints are denoted as a triple ($h,r,t$) where $h,r$ and $t$ are a head entity, a relation, and a tail entity respectively. Since KGs can represent complex relations between entities, they serve as key components for knowledge-intensive applications Ji et al. (2021); Xiong et al. (2017); He et al. (2017); Zhang et al. (2016); Liu et al. (2021).
Despite their applicability, real-world KGs miss factual relations, which can be inferred from existing facts in the KGs. Hence, the task of knowledge graph completion (KGC) has become an active research topic Ji et al. (2021). Given an incomplete triple $(h,r,?)$, this task is to predict the correct tail $t$. A true triple $(h,r,t)$ in KG and a false triple $(h,r,\hat{t})$ which does not exist in KG are called positive and negative, respectively. A negative triple difficult for a KGC method to distinguish from its corresponding positive triple is regarded as a hard negative triple.
Existing KGC methods are categorized into two approaches. An embedding-based approach learns embeddings of entities in continuous vector spaces, but ignores contextualized text information in KGs, thus being inapplicable to entities and relations unseen in training (Bordes et al., 2013; Balazevic et al., 2019). A text-based approach, based on pretrained language models (PLMs), learns textual representations of KGs, but suffers from a lack of structural knowledge in KGs (Wang et al., 2022a).
Meanwhile, contrastive learning has become a key component of representation learning Kalantidis et al. (2020); Robinson et al. (2021); Gao et al. (2021); Wang et al. (2022b), but an important aspect of contrastive learning, i.e., the effect of hard negatives, has so far been underexplored in KGC. In this paper, we empirically validate significant relationships between the structural properties of KGs and the performance of the PLM-based KGC methods.
To address the aforementioned longstanding limitations of two KGC approaches by utilizing the above relationships, we hypothesize that incorporating the structural inductive bias of KGs into both sampling hard negatives and fine-tuning PLMs leads to a major breakthrough in learning comprehensive representations of KGs. To this end, we propose a Subgraph-Aware Training framework for KGC (SATKGC), which (i) samples subgraphs of a KG to treat triples of each subgraph as a mini-batch to encourage hard negative sampling, and (ii) fine-tunes a PLM via contrastive learning which focuses more on structurally harder entities and structurally harder negative triples induced by topological bias in the KG. To sum up, we make four contributions.
  • We provide key insights that the topological structure of KGs is closely related to the performance of PLM-based KGC methods.
  • We propose a subgraph-aware training strategy for PLM-based KGC methods, which is effective in sampling in-batch hard negatives.
  • We propose a novel contrastive learning method that gives different importances to positive and negative triples based on the structural properties of KGs.
  • We conduct extensive experiments on four KGC benchmarks to demonstrate the superiority of SATKGC over existing KGC methods.
2 Related Work
An embedding-based approach maps complex and structured knowledge into low-dimensional spaces. This approach computes the plausibility of a triple using translational scoring functions on the embeddings of the triple’s head, relation, and tail Bordes et al. (2013); Sun et al. (2019); Zhang et al. (2022); Ge et al. (2023), e.g., $h+r\approx t$, or semantic matching functions which match latent semantics of entities and relations Nickel et al. (2011); Yang et al. (2015); Trouillon et al. (2016); Balazevic et al. (2019). This approach exploits the spatial relations of the embeddings, but cannot make use of texts in KGs, i.e., the source of semantic relations.
In contrast, a text-based approach learns contextualized representations of the textual contents (e.g., names and descriptions) of entities and relations by leveraging PLMs Yao et al. (2019); Wang et al. (2021b); Xie et al. (2016); Daza et al. (2021); Kim et al. (2020); Wang et al. (2022a); Chen et al. (2022); Zhang et al. (2023a). More recently, there has been a significant increase in the adoption of large language models (LLMs) in KGC Zhang et al. (2023c). However, these PLM-based models are often oblivious to structural inductive bias in KGs.
A few attempts have been made to utilize the above two approaches at once. StAR Wang et al. (2021d) proposes an ensemble model incorporating an output of a Transformer encoder Vaswani et al. (2017) with a triple score produced by RotatE Sun et al. (2019). CSProm-KG Chen et al. (2023) trains KG embeddings through the soft prompt for a PLM. Nevertheless, the integration of structural and textual information in a KG in training has not yet been fully realized.
Contrastive learning, shown to be effective in various fields Wu et al. (2018); Haque et al. (2022); Fang et al. (2020); Zhang et al. (2023b); Gao et al. (2021), has recently emerged as a promising approach in the context of KGC Wang et al. (2022a); Yang et al. (2020); Qiao et al. (2023). Despite a critical aspect of contrastive learning, the effect of hard negatives has been overlooked in KGC.
Random walk with restart (RWR) (Page et al., 1999) and its extension, biased random walk with restart (BRWR), have been employed in various domains such as node representation learning (Perozzi et al., 2014) and graph traversals (Ahrabian et al., 2020). In BRWR, a random walker performs random walks in a graph from the source node. For each iteration, the walker moves from the current node $u$ to either (i) source with a probability of $p_{r}$, or (ii) one of the neighbors of $u$ with a probability of ${1-p_{r}}$, where $p_{r}$ is a hyperparameter. In case (ii), the probability of selecting one of the neighbors is decided by a domain-specific probability distribution, whereas one is selected uniformly at random in RWR. To our knowledge, we are the first to extract a subgraph of KG via BRWR to utilize the subgraph as a mini-batch during training.
3 Motivation
To demonstrate the limitations of text-based methods exhibiting competitive performance such as SimKGC (Wang et al., 2022a) and StAR (Wang et al., 2021a), we investigate the characteristics of false positive (FP) triples which are ranked by these models higher than corresponding true triples, on two widely-used datasets WN18RR and FB15k-237. Our analysis draws two conclusions.
Figure 1: False positive (FP) ratio against the distance (i.e., length of the shortest path) between head and tail of a FP triple in KG across different text-based methods.
Figure 2: False positive (FP) ratio against the degree of head for a FP triple across different text-based methods.
First, the closer the tail and head of a false triple are to each other in the KG, the more likely the false triple is to be ranked higher than the corresponding true triple. Figure 1 illustrates the distribution of distance, i.e., the length of the shortest path, between the head and tail of a FP triple, where $y$-axis represents the FP ratio111(the number of FPs with a specific head-to-tail distance) / (the number of pairs of entities in the KG with the same distance). for each distance. For StAR and SimKGC, the FP ratio dramatically grows as the distance decreases (see green and red bars in Figure 1). These findings highlight the importance of considering the proximity of two entities of a negative triple in KGs for text-based methods to distinguish a positive triple from the negative.
Figure 3: Overview of the proposed training framework, which consists of: (i) sampling subgraphs from KG (before training); (ii) selecting a subgraph; (iii) fetching a mini-batch of triples in the subgraph; (iv) calculating a similarity between embeddings of every (head, relation) pair and every tail in the mini-batch; (v) contrastive learning via infoNCE loss incorporated with two structure-aware factors.
Second, we discover that the higher the degree of the head in a false triple is, the more likely the false triple is to be ranked higher than the corresponding true triple. Figure 2 illustrates the distribution of the degree of heads of FP triples. The FPs are sorted in ascending order of the degrees, and then they are divided into five groups such that each group contains an equal number of distinct degrees. The $y$-axis represents the FP ratio222(the average number of FPs whose head’s degree falls into each group) / (the number of entities in the KG whose degree falls into each group) in each degree group. The FP ratio for StAR and SimKGC increases as the degree of the head grows (see green and red bars in Figure 2). This indicates that the existing text-based methods have difficulty in predicting correct tails for missing triples with the high-degree heads. Hence, the degree can be taken into account to enhance the performance of text-based methods.
Our proposed framework (dubbed SATKGC) tackles the above two phenomena [^3] The trends in Figures 1 and 2 are also confirmed on Wikidata5M and NELL-995., thereby significantly reducing the FPs for all distances and all degree groups compared to the existing methods (see blue bars in Figures 1 and 2).
4 Method
We propose a novel training framework for KGC that captures the structural inductive bias of the KG, based on the aforementioned observations.
Figure 3 illustrates the overview of our framework. First, for every triple, a subgraph is extracted around that triple from the KG, performed before training (Section 4.1). During training, we keep track of the number of visits for every triple. For each iteration, a subgraph is selected based on that number, and then all forward and inverse triples in the subgraph are fetched as a mini-batch $B$ to a model (Section 4.2). We adopt the bi-encoder architecture Wang et al. (2022b) as a backbone, which uses pre-trained BERT (Devlin et al., 2018) as encoders. Specifically, $\textit{Encoder}_{hr}$ and $\textit{Encoder}_{t}$ take the text, i.e., name and description, of $(h,r)$ and $t$ as input, and produce their embeddings $x_{hr}$ and $x_{t}$ respectively. We then calculate the cosine similarity between $x_{hr}$ and $x_{t}$ for every $(h,r)$ and $t$ in the mini-batch, and perform contrastive learning based on two structure-aware factors (Section 4.3). The model inference is described in Appendix A.
Figure 4: Example of BRWR-based subgraph sampling; (a) probability of selecting the start entity $s$ between $h$ and $t$ of a center triple, where $t$ with a lower degree is more likely to be selected as $s$ than $h$; (b) probability of selecting a neighbor of current entity $u$. A random walker is more likely to move to $v_{1}$ than to $v_{2}$ whose degree is larger than $v_{1}$.
4.1 Random-Walk Based Subgraph Sampling
First, we aim to extract subgraphs from the KG to treat all the triples in the extracted subgraph as a mini-batch for training. For each triple in the KG, we perform BRWR starting from that triple called a center triple, and the triples visited by BRWR composes an extracted subgraph as follows: (i) for each center triple, either head $h$ or tail $t$ of the center triple is selected as the start entity $s$ based on an inverse degree distribution of $h$ and $t$, i.e., $\frac{|N(x)|^{-1}}{|N(h)|^{-1}+|N(t)|^{-1}}$, where $x\in\{h,t\}$ and $N(x)$ denotes a set of $x$’s neighbors; (ii) next, we perform BRWR from $s$ until we sample $M$ triples where $M$ is a predefined maximum number (e.g., 10,000). For each iteration in BRWR, a random walker moves from the current node to either $s$ with a probability of $p_{r}$ or one of the neighbors of the current node with a probability of $1-p_{r}$. We define the probability of selecting one of $u$’s neighbors $v\in N(u)$ as $p_{v}=\frac{|N(v)|^{-1}}{\sum_{v\in N(u)}|N(v)|^{-1}}$, which is a normalized inverse degree distribution of the neighbors. Figures 4(a) and 4(b) show the running example of step (i) and an iteration of step (ii).
Performed before the model training, this subgraph sampling algorithm can extract many distinct entities due to the inverse degree distribution, which will be validated in Section 5.5.
4.2 Subgraph as a Mini-Batch
In this subsection, we describe our training framework dubbed Subgraph as a Mini-batch (SaaM). We count the number of visits for all triples in the training set $T$ throughout the training process. For every iteration, we select the subgraph whose center triple is the least frequently visited, to prioritize unpopular and peripheral triples. The rationale behind this selection will be elaborated in Section 6.2. Next, we randomly select $|B|/2$ triples from the subgraph, and feed to the model a mini-batch $B$ of the these selected triples $(h,r,t)$ and their inverse triples $(t,r^{-1},h)$. For every positive triple $(h,r,t)\in B$, we obtain negative triples $(h,r,\hat{t})$ with $t$ replaced by $|B|-1$ tails $\hat{t}$ of the other triples in $B$. As per our observation in Figure 1, these negative triples are likely to be hard negatives, which will facilitate contrastive learning.
To introduce a concept corresponding to an epoch in our training framework, let iterating above process for $|T|/|B|$ times, i.e., selecting and feeding $|T|/|B|$ subgraphs, be a “phase” as opposed to an epoch that visits every triple exactly once.
4.3 Subgraph-Aware Contrastive Learning
For effective contrastive learning, we propose to incorporate two structure-aware factors into InfoNCE loss with additive margin (Chen et al., 2020; Yang et al., 2019) over a mini-batch: (a) structural hardness of negatives for each positive triple, and (b) structural hardness of head entities in the mini-batch. These two factors are simply obtained from the topological structure of the subgraph, while incurring a minimal computational overhead overall.
4.3.1 Structural Hardness of Negative Triples
Entities close to each other are more likely to be related than entities far away from each other. Although text-based KGC methods capture semantic relations within the text of triples, they overlook capturing the proximity between two entities in a negative triple in KGs. To tackle this problem, we propose a novel InfoNCE loss that penalizes structurally hard negative triples. For each positive triple $(h,r,t)$ in a mini-batch $B$ of SaaM, the loss function $L_{(h,r,t)}$ is defined as
$L_{(h,r,t)}=-\log\frac{e^{(\phi(h,r,t)-\gamma)/\tau}}{e^{(\phi(h,r,t)-\gamma)/\tau}+\sum\limits_{i=1}^{|B|-1}e^{(\phi(h,r,t_{i})+\beta\omega_{ht_{i}})/\tau}}\end{aligned}$
where $\gamma$ is an additive margin, a temperature parameter $\tau$ adjusts the importance of negatives, structural hardness $\omega_{ht_{i}}$ stands for how hard a negative triple ($h,r,t_{i}$) is in terms of the structural relation between $h$ and $t_{i}$ in the KG, and $\beta$ is a trainable parameter that adjusts the relative importance of $\omega_{ht_{i}}$. Specifically, we define $\omega_{ht_{i}}$ as the reciprocal of the distance (i.e., length of the shortest path) between $h$ and $t_{i}$ to impose a larger weight $\omega_{ht_{i}}$ to the negative triple with a shorter distance between $h$ and $t_{i}$, which serves as a harder negative triple.
Since computing the exact distance between every head and every tail in $B$ may spend considerable time, in our implementation, we calculate the approximate distance between $h$ and $t_{i}$, i.e., the multiplication of two distances: (d1) the distance between $h$ and head $h_{c}$ of the center triple of $B$, and (d2) the distance between $t$ and $h_{c}$. For this, the distance between $h_{c}$ and every entity in $B$ is pre-computed before training, the multiplication between the two distances is performed in parallel during training, which requires a minimal computational overhead. [^4] We conducted experiments using the exact distance and different approximate distances, e.g., the sum of (d1) and (d2), but the performance gap between all the methods is very marginal.
4.3.2 Structural Hardness of Entities
For many triples with the same head in the KG, varying only relation $r$ in $(h,r,?)$ may lead to different correct tails with various semantic contexts. The text-based KGC methods may find it difficult to predict the correct tail for these triples, as their head-relation encoders may generate less diverse embeddings for $(h,r)$ due to much shorter text of relation $r$ than entity $h$.
To encourage the text-based method to more sensitively adapt to varying relations for many triples with the identical head, we propose a novel loss weighting strategy that penalizes structurally hard head entities. For each triple in a mini-batch $B$ in SaaM, mini-batch loss $L_{B}$ is defined as
$L_{B}=\sum_{(h,r,t)\in B}\psi_{h}L_{(h,r,t)}$
where structural hardness $\psi_{h}$ of head $h$ indicates how difficult $(h,r,t)$ is in terms of the structural property of $h$ in the KG. Specifically, we define $\psi_{h}$ as $log(d_{h}+1)$ where $d_{h}$ represents the degree of $h$, and adding one to $d_{h}$ can prevent $\psi_{h}$ from becoming zero when $d_{h}=1$. To sum up, $\psi_{h}$ ensures that the triples whose heads have a larger degree contribute more significantly to $L_{B}$.
dataset #entity #relation #train #valid #test
WN18RR 40,943 11 86,835 3,034 3,134
NELL-995 74,332 200 149,678 543 3,992
FB15k-237 14,541 237 272,115 17,535 20,466
Wikidata5M-Trans 4,594,485 822 20,614,279 5,163 5,163
Wikidata5M-Ind 4,579,609 822 20,496,514 6,699 6,894
Table 1: Statistics of datasets.
Approach WN18RR NELL-995 FB15k-237
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
Embedding-based approach
TuckER Balazevic et al. (2019) 0.466 0.432 0.478 0.518 0.423 0.363 0.455 0.536 0.361 0.265 0.391 0.538
RotatE Sun et al. (2019) 0.471 0.421 0.490 0.568 0.411 0.350 0.439 0.522 0.335 0.243 0.374 0.529
KGTuner Zhang et al. (2022) 0.481 0.438 0.499 0.556 0.428 0.371 0.458 0.544 0.345 0.252 0.381 0.534
CompoundE Ge et al. (2023) 0.492 0.452 0.510 0.570 0.434 0.381 0.466 0.550 0.350 0.262 0.390 0.547
CSProm-KG Chen et al. (2023) 0.569 0.520 0.590 0.675 0.422 0.381 0.474 0.535 0.355 0.261 0.389 0.531
Text-based approach
StAR Wang et al. (2021a) 0.398 0.238 0.487 0.698 0.419 0.238 0.395 0.433 0.288 0.195 0.313 0.480
HaSa Zhang et al. (2023a) 0.538 0.444 0.588 0.713 0.411 0.310 0.431 0.524 0.304 0.220 0.325 0.483
KG-S2S Chen et al. (2022) 0.572 0.529 0.595 0.663 0.416 0.311 0.421 0.519 0.337 0.255 0.374 0.496
SimKGC Wang et al. (2022a) 0.671 0.580 0.729 0.811 0.425 0.353 0.447 0.531 0.340 0.252 0.365 0.515
GHN Qiao et al. (2023) 0.678 0.596 0.719 0.821 - - - - 0.339 0.251 0.364 0.518
Ensemble approach
StAR Wang et al. (2021a) 0.520 0.456 0.509 0.707 0.415 0.311 0.402 0.510 0.332 0.229 0.387 0.526
SATKGC-SPW-DW 0.672 0.605 0.714 0.811 0.411 0.310 0.431 0.524 0.345 0.225 0.367 0.520
SATKGC-DW 0.679 0.609 0.723 0.813 0.429 0.361 0.466 0.551 0.351 0.261 0.372 0.525
SATKGC-SPW 0.681 0.612 0.729 0.820 0.431 0.369 0.472 0.556 0.359 0.271 0.393 0.538
SATKGC 0.689 0.621 0.731 0.823 0.433 0.389 0.476 0.560 0.368 0.276 0.401 0.548
Table 2: KGC results for the WN18RR, NELL-995, and FB15k-237 datasets. “SPW” and “DW” refer to shortest path weight and degree weight respectively. The best and second-best performances are denoted in bold and underlined respectively. $\dagger$: numbers are from Qiao et al. (2023).
Approach Wikidata5M-Trans Wikidata5M-Ind
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
Embedding-based approach
TransE Bordes et al. (2013) 0.253 0.170 0.311 0.392 - - - -
RotatE Sun et al. (2019) 0.290 0.234 0.322 0.390 - - - -
Text-based approach
DKRL Xie et al. (2016) 0.160 0.120 0.181 0.229 0.231 0.059 0.320 0.546
KEPLER Wang et al. (2021c) 0.212 0.175 0.221 0.276 0.403 0.225 0.517 0.725
BLP-ComplEx Daza et al. (2021) - - - - 0.491 0.261 0.670 0.881
BLP-SimplE Daza et al. (2021) - - - - 0.490 0.283 0.641 0.868
SimKGC Wang et al. (2022a) 0.358 0.313 0.376 0.441 0.714 0.609 0.785 0.917
SATKGC 0.404 0.361 0.423 0.479 0.763 0.659 0.809 0.927
Table 3: KGC results for Wikidata5M-Trans (transductive setting) and Wikidata5M-Ind (inductive setting). The results for embedding-based approach on Wikidata5M-ind are missing as they cannot be used in the inductive setting. Additionally, BLP-ComplEx Daza et al. (2021) and BLP-SimplE Daza et al. (2021) results on Wikidata5M-Trans are missing because they are inherently targeted for inductive KGC.
5 Experiments
5.1 Experimental Setup
FB15K-237N
Models MRR Hits@1 Hits@3 Hits@10
KoPA (Zhang et al., 2023c) 0.483 0.344 0.559 0.721
SATKGC 0.668 0.557 0.742 0.858
Table 4: KGC results on FB15K-237N. A correct tail is ranked among 1,000 randomly-selected entities due to the long inference time of KoPA.
For evaluation we adopt widely-used KG datasets WN18RR, FB15k-237, NELL-995 and Wikidata5M. Table 1 shows their statistics, and more details are provided in Appendix C.
For every incomplete triple in the test set, we compute mean reciprocal rank (MRR) and Hits@$k$ where $k\in\{1,3,10\}$ as evaluation metrics based on the rank of the correct entity among all the entities in the KG. We use the mean of the forward and backward prediction results as the final performance measure. Implementation details are described in Appendix D.
5.2 Main Results
We compare SATKGC with existing embedding-based and text-based approaches. Table 2 shows the results on WN18RR, NELL-995, and FB15k-237, and Table 3 shows the results on Wikidata5M-Trans and Wikidata5M-Ind. [^5] The results of GHN on NELL-995 are unavailable due to a lack of access to the requisite codes. SATKGC denotes the bi-encoder architectures trained by our learning framework in Figure 3. Note that Table 2 presents the performance of the latest baselines, and the complete results for all baselines compared with ours are found in Appendix K. SATKGC consistently outperforms all the existing methods on all the datasets except MRR on NELL-995. SATKGC demonstrates significantly higher MRR and Hits@1 than other baselines especially on WN18RR and FB15k-237. The small performance improvement of our method on NELL-995 stems from the limited textual context in this dataset, unlike other datasets. Nevertheless, SATKGC is the runner-up lagged only 0.001 behind CompoundE for MRR on NELL-995, while outperforming all baselines on the remaining metrics.
As shown in Table 3, SATKGC demonstrates its applicability to large-scale KGs, and achieves strong performance in both inductive and transductive settings. [^6] Baselines listed in Table 2 but not in Table 3 could not be evaluated on Wikidata5M due to out-of-memory for StAR and CSProm-KG, or unreasonably large training time, i.e., more than 100 hours expected, for the remaining baselines. SATKGC significantly outperforms the baselines on Wikidata5M-Ind, reaching MRR of 0.763 and Hits@1 of 0.659.777Wikidata5M-Ind shows better performance than Wikidata5M-Trans, because a model ranks 7,475 entities in the test set for Wikidata5M-Ind while ranking about 4.6 million entities for Wikidata5M-Trans. In the transductive setting, performance degrades in the order of WN18RR, Wikidata5M-Trans, and FB15k-237, showing that a higher average degree of entities tends to negatively affect performance.
To compare our framework employing BERTs with a LLM-based model, we evaluate the KGC performance of KoPA (Zhang et al., 2023c), which adopts Alpaca (Taori et al., 2023) fine-tuned with LoRA (Hu et al., 2021) as its backbone, on the FB15k-237N dataset (Zhang et al., 2023c). Since KoPA cannot perform inference for all queries within a reasonable time, i.e., 111 hours expected, we rank the correct tail among 1,000 randomly selected entities for both KoPA and SATKGC. As shown in Table 4, SATKGC outperforms KoPA on all metrics, demonstrating that LLMs do not necessarily produce superior results on KGC.
5.3 Ablation study
To show the contributions of applying SaaM and adding the two weighting factors to InfoNCE loss, we compare results of four different settings including SATKGC in Table 2: SATKGC-SPW-DW denotes applying only SaaM and using original InfoNCE loss, SATKGC-SPW represents applying SaaM and the degree factor $\psi_{h}$, and SATKGC-DW represents applying SaaM and the distance factor $\omega_{ht_{i}}$. SATKGC-SPW-DW already shows higher Hits@1 than other baselines on WN18RR, which highlights that SaaM alone can lead to performance improvement. The efficacy of SaaM is further evidenced in Appendix F. Between SATKGC-SPW and SATKGC-DW, SATKGC-SPW achieves higher performance, which indicates that the degree factor contributes more than the distance factor.
WN18RR
Encoders MRR Hits@1 Hits@3 Hits@10 Parameters
BERT-base 0.689 0.621 0.731 0.823 220M
DeBERTa-base 0.689 0.631 0.736 0.832 280M
BERT-large 0.685 0.619 0.723 0.817 680M
DeBERTa-large 0.706 0.638 0.759 0.854 800M
NELL-995
Encoders MRR Hits@1 Hits@3 Hits@10 Parameters
BERT-base 0.433 0.389 0.476 0.560 220M
DeBERTa-base 0.434 0.388 0.478 0.559 280M
BERT-large 0.432 0.367 0.475 0.541 680M
DeBERTa-large 0.426 0.368 0.428 0.551 800M
Table 5: Performance comparison for different encoders.
5.4 Performance Across Encoders
To investigate the impact of the encoder architecture and the number of model parameters, we conduct experiments replacing BERT-base in SATKGC with BERT-large, DeBERTa-base, and DeBERTa-large (He et al., 2020). Table 5 presents the results. SATKGC is highly compatible with different encoders, showing the competitive performance. DeBERTa-large fine-tuned by SATKGC achieves the best performance on WN18RR. In addition, an increase in the number of model parameters may not necessarily result in enhanced performance on KGC, e.g., BERT-large on WN18RR, and BERT-large and DeBERTa-large on NELL-995 underperform the smaller encoders.
5.5 Comparing Sampling Methods
WN18RR
Methods MRR Hits@1 Hits@3 Hits@10
RWR 0.682 0.620 0.728 0.821
BRWR 0.689 0.621 0.731 0.823
BRWR_P 0.673 0.610 0.726 0.810
MCMC 0.676 0.609 0.716 0.802
FB15k-237
Methods MRR Hits@1 Hits@3 Hits@10
RWR 0.355 0.264 0.382 0.532
BRWR 0.368 0.279 0.403 0.548
BRWR_P 0.347 0.258 0.376 0.526
MCMC 0.346 0.259 0.376 0.521
Table 6: Experiments comparing subgraph sampling methods on WN18RR and FB15k-237.
We investigate how model performance varies depending on the probability distribution $p_{v}$ used for neighbor selection in Section 4.1.888Recall that in our BRWR algorithm, a random walker selects one of the neighbors $v$ of a current node based on the inverse degree distribution $p_{v}$. We compare the performance of SATKGC using $p_{v}$ in subgraph sampling with two variants, one with $p_{v}$ replaced by the uniform distribution (dubbed RWR) and the other with $p_{v}$ replaced by the degree proportional distribution (dubbed BRWR_P). Table 6 shows the results. The three methods mostly outperform existing KGC models in Hits@1, with BRWR performing best. BRWR_P performs the worst, likely due to many duplicate entities in the extracted subgraphs sampled from the degree-proportional distribution. We also employ a Markov chain Monte Carlo (MCMC) based subgraph sampling method (Yang et al., 2020), referred to as MCMC (see details of MCMC in Appendix B). Note that in Table 2, MCMC outperforms other methods in terms of Hits@1 on WN18RR.
6 Analysis
6.1 Analysis on Negative Triples
Figure 5: Percentage of in-batch negative triples on FB15k-237 according to the range of cosine similarity scores predicted by SimKGC for different epochs and by SATKGC for different phases.
Figure 5 shows how the cosine similarity distribution of in-batch negative triples varies depending on the epoch of SimKGC, and the phase of SATKGC for FB15k-237. SATKGC encounters consistently more hard negatives with scores from 0.2 to 1.0 than SimKGC, though the majority of the scores range from -0.2 to 0.2 by the end of training for both methods. [^9] This trend is also observed in WN18RR. We speculate that SATKGC ends up with distinguishing positives from the hard negatives sampled from the subgraphs of a KG, as opposed to SimKGC which employs randomly-sampled easy negatives. The effectiveness using only negatives from SaaM in enhancing performance is demonstrated in Appendix G.
Based on our analysis, only 2.83% and 4.42% of the true triples ranked within the top 10 by SimKGC drop out of the top 10 for SATKGC on WN18RR and FB15k-237, respectively. In contrast, 34.87% and 13.03% of the true triples dropping out of the top 10 for SimKGC are ranked within the top 10 by SATKGC. This indicates that SATKGC effectively samples hard negatives while reducing false negatives.
6.2 Skewness of Occurrences
Figure 6: (a) Number of occurrences of triples; (b) number of occurrences of entities. Both are counted throughout the entire training process for RANDOM and the SaaM variants, i.e., RWR, BRWR_P, and BRWR.
Figures 6(a) and 6(b) show the number of visits of every triple and every entity, respectively, in the training process for RANDOM and the SaaM variants: RWR, BRWR_P, and BRWR. RANDOM, adopted in all the baselines, denotes selecting a mini-batch of triples at random without replacement. Triples and entities are sorted by descending frequency. Figure 6(a) demonstrates that RANDOM exhibits a uniform distribution, while RWR, BRWR_P, and BRWR display varying degrees of skewness, with BRWR being the most skewed and BRWR_P the least. Figure 6(b) illustrates that BRWR is the least skewed whereas RANDOM shows the most skewed distribution. As a result, a larger skewness in the distribution of number of visits for triples in turn leads to more equally visiting entities, thus improving the performance. [^10] The similar trends are shown on WN18RR.
Further analysis reinforces this finding. In FB15k-237, the average degree of FP triples’ tails is 75 for SATKGC and 63 for SimKGC. A smaller portion of low-degree tails for SATKGC than for SimKGC indicates that exposure to more low-degree entities $\hat{t}$ in training helps the model position their embeddings farther from the $(h,r)$ embeddings for negative triples $(h,r,\hat{t})$, as SATKGC visits low-degree entities more often during training than RANDOM for SimKGC. [^11] This is also confirmed in all the other datasets.
For BRWR in Figure 6(a), we examine the structural characteristics on sets $S_{m}$ and $S_{l}$ of entities in $1,000$ most and least frequent triples, respectively. The entities in $S_{m}$ have an average degree of $11.1$, compared to $297.3$ for those in $S_{l}$. The betweenness centrality [^12] The betweenness centrality of node $v$ is the number of shortest paths that pass through $v$ in the graph divided by the total number of shortest paths between all pairs of nodes, which measures $v$’s importance in the graph. averages around $5.2\times 10^{-5}$ for $S_{m}$ and $8.2\times 10^{-4}$ for $S_{l}$. These observations implies that SaaM prioritizes visiting unpopular and peripheral triples in the KG over focusing on information-rich triples.
7 Conclusion
In this paper, we propose a generic training scheme and a new contrastive learning method for KGC. Harmonizing (i) SaaM using a subgraph of KG as a mini-batch, and (ii) contrastive learning that incorporates the structural hardness of the KG into fine-tuning PLMs helps learning the contextual text embeddings aware of the difficulty in the structural context of the KG. Our findings imply that unequally feeding triples in training and leveraging the unique characteristics of KG lead to the effective text-based KGC method, achieving state-of-the-art performance on the four KG benchmarks.
Limitations
While our training scheme efficiently samples many hard negatives from subgraphs, BRWR incurs a computational overhead to extract subgraphs from Wikidata5M before training (see Appendix E). However, note that this algorithm can be efficiently parallelized or replaced by a simpler one.
Appendix A Inference
For inference, the bi-encoder model calculates the cosine similarity between $\textbf{x}_{hr}$ for a given $(h,r,?)$ and $\textbf{x}_{t}$ for all entities $t$. Then the tails with the top-k largest cosine similarities are answers. For a single pair, we need $|E|$ forward passes of $\textit{Encoder}_{t}$ to obtain $\textbf{x}_{t}$ for all entities $t$. Given a set $T$ of test triples, $2|T|$ forward passes of $\textit{Encoder}_{hr}$ are required to get $\textbf{x}_{hr}$ for every triple $(h,r,?)\in T$ and its inverse triple $(t,r^{-1},?)$, thus resulting in $O(|E|+|T|)$ computation in total.
Appendix B Markov Chain Monte Carlo Based Subgraph Sampling
Input: DFS path $D=\{t_{1},t_{2},...,t_{d}\}$, proposal distribution $q$, number $k$ of negative samples, $burn\_in$ period
Output: negative tails
Initialize current negative node $x$ at random
$i\leftarrow 0$, $S\leftarrow\emptyset$
for each triple $t$ in $D$ do
       Initialize $j$ as 0
       $h,r\leftarrow$ head, relation in triple
       if $i\leq burn\_in$ then
             $i\leftarrow i+1$
             Sample an entity $y$ from $q(y|x)$
             Generate $r\in\left[0,1\right]$ uniformly at random
             if $r\leq\min(1,\frac{cos(\textbf{x}_{hr},\textbf{x}_{y})^{\alpha}}{cos(\textbf{x}_{hr},\textbf{x}_{x})^{\alpha}}\frac{q(x|y)}{q(y|x)})$ then
                   $x\leftarrow y$
      else
             while $j<k$ do
                   Sample an entity $y$ from $q(y|x)$
                   Add $y$ to $S$
                   $x\leftarrow y$
                   $j\leftarrow j+1$
return $S$;
Algorithm 1 MCMC-based Subgraph Sampling
Inspired by a negative sampling approach (Yang et al., 2020) based on Markov chain Monte Carlo (MCMC) in a general graph, we propose a new method to sample subgraphs from a KG. A negative sampling distribution should be positively but sublinearly correlated with the positive sampling distribution, which was validated by Yang et al. (2020). To include the entities close to a positive triple in the KG in negative triples, we define the sampling distribution $p_{n}$ of the negative tail $\hat{t}$ as : $p_{n}(\hat{t}|h,r)\propto p_{d}(\hat{t}|h,r)^{\alpha},0<\alpha<1$, $p_{d}(\hat{t}|h,r)=\frac{\cos(\textbf{x}_{hr},\textbf{x}_{\hat{t}})}{\sum_{e\in E}\cos(\textbf{x}_{hr},\textbf{x}_{e})}$. where $\alpha$ is a parameter to stabilize the optimization process, $E$ is a set of entities in the KG, and $p_{d}$ is the sampling distribution of the positive tail. Calculating the normalization term in $p_{d}(\hat{t}|h,r)$ is time consuming and almost impossible. Therefore, we sample a negative tail $\hat{t}$ from $\tilde{p_{d}}=\cos(\textbf{x}_{hr},\textbf{x}_{\hat{t}})$ by using the Metropolis-Hastings (M-H) algorithm, and randomly select a triple whose head is $\hat{t}$.
Algorithm 1 describes the process of sampling a subgraph by the M-H algorithm. To prepare a mini-batch of triples for SaaM, we traverse KG using depth-first search (DFS). From each entity $e\in E$, we perform DFS until we visit $d$ triples. For every visited triple along the DFS path, the triple inherits the probability distribution $\tilde{p_{d}}$ from the previous triple in the path, and $k$ negative tails $\hat{t}$ are sampled from this distribution. Then $\tilde{p_{d}}$ is updated. The proposal distribution $q$ is defined as a mixture of uniform sampling and sampling from the nearest $k$ nodes with the 0.5 probability each (Yang et al., 2020). Both $d$ and $k$ above are hyperparameters. The $d$ triples in a DFS path, the sampled $d\times k$ triples, and their inverse triples compose a subgraph. We throw away the tails $\hat{t}$ extracted during the burn-in period, and use the tails extracted after the period as the heads of the triples in the subgraph.
Appendix C Datasets
In this paper, we use four KGC benchmarks. WN18RR is a sparse KG with a total of 11 relations and $\sim 41k$ entities. WN18RR is the dataset derived from WN18, consisting of relations and entities from WordNet Miller (1992). WN18RR addresses the drawbacks of test set leakage by removing the inverse relation in WN18. NELL-995 is a sparse KG extracted from the web. FB15k-237 is a dense KG with 237 relations. Wikidata5M, a much larger KG than the others, provides transductive and inductive settings. Wikidata5M-Trans is for the transductive setting, where entities are shared and triples are disjoint across training, validation, and test. Wikidata5M-Ind is for the inductive setting, where the entities and triples are mutually disjoint across training, validation, and test (Wang et al., 2021c).
Appendix D Implementation Details
In our weighted InfoNCE loss, additive margin $\gamma$ is set to 0.02. We select the best performing batch sizes of 1024 from 512, 1024, 1536, 2048 for WN18RR, NELL-995, and Wikidata5M, and 3072 from 1024, 2048, 3072, 4096 for FB15k-237. We set the restart probability $p_{r}$ to 1/25 in BRWR. We used six A6000 GPUs and 256G RAM. Training on WN18RR took 50 phases, for a total of 4 hours. NELL-995 and FB15k-237 took 20 phases and a total of 3 hours and 10 hours respectively, while Wikidata5M took 2 phases and 20 hours.
WN18RR
Size MRR Hits@1 Hits@3 Hits@10
512 0.591 0.571 0.604 0.713
1024 0.689 0.617 0.741 0.831
1536 0.679 0.608 0.732 0.818
2048 0.652 0.598 0.668 0.715
FB15k-237
Size MRR Hits@1 Hits@3 Hits@10
1024 0.322 0.234 0.347 0.500
2048 0.339 0.250 0.368 0.521
3072 0.368 0.279 0.400 0.548
4096 0.357 0.261 0.357 0.532
Table 7: The results of investigating the model performance with respect to the batch size.
Appendix E Runtime Analysis
Sampling Time
Algorithm WN18RR FB15k-237 Wiki5M-Trans Wiki5M-Ind
BRWR 13m 1h 14h 14h
Shortest Path + Degree 7m 15m 7h 7h
Training Time (SATKGC: time per phase, others: time per epoch)
Model WN18RR FB15k-237 Wiki5M-Trans Wiki5M-Ind
SATKGC 7m 30m 10h 10h
SimKGC Wang et al. (2022a) 4m 20m 9h 9h
StAR Wang et al. (2021a) 1h 1h 30m - -
SANS Ahrabian et al. (2020) 45m 57m - -
Table 8: The elapsed time required for sampling and the time allotted per phase during training.
Training SATKGC incurs a marginal computational overhead because (i) sampling subgraphs and (ii) computing distances and degrees are performed in advance before training. As shown in Table 8, the computational cost for (i) and (ii) is acceptable. The cost depends on the size of a mini-batch, which can be adjusted. Moreover, the time complexity for (ii) is acceptable. For each mini-batch $B$ of triples, we run Dijkstra’s single source shortest path algorithm, and thus the runtime to compute the distance is O($|V|log|V|+|B|log|V|$), where $V$ is a set of entities in $B$. As described in Section 4.3.1, we do not calculate all-pair shortest paths for every pair of vertices in $V$. To reduce the computational overhead, we compute the approximate distance by using the shortest path between the head of the center triple in a subgraph and every tail in that subgraph, which have been already obtained fro Dijkstra’s algorithm. Table 8 shows the training time per phase for SATKGC and per epoch for SimKGC Wang et al. (2022a), StAR Wang et al. (2021a), and SANS Ahrabian et al. (2020). SATKGC remains competitive, though it takes slightly more time than SimKGC due to the computational overhead for (a) computing its loss using the shortest path weight (SPW) and degree weight (DW), (b) counting the occurrences of visited triples to select the next subgraph, and (c) fetching subgraphs. StAR and SANS take longer than SATKGC. StAR runs out of memory on the Wikidata5M datasets, while SANS cannot be applied to the inductive setting (Wikidata5M-Ind), and is not expected to finish within a reasonable time on Wikidata5M-Trans.
WN18RR
Method MRR Hits@1 Hits@3 Hits@10
StARWang et al. (2021a) 0.398 0.238 0.487 0.698
StAR+SaaM 0.411 0.261 0.511 0.729
FB15k-237
Method MRR Hits@1 Hits@3 Hits@10
StARWang et al. (2021a) 0.288 0.195 0.313 0.480
StAR+SaaM 0.319 0.220 0.334 0.490
Table 9: Performance comparison between original StAR and StAR+SaaM where StAR+SaaM stands for the StAR model architecture trained by our training framework SaaM.
Appendix F Effectiveness of SaaM
To demonstrate the generality of SaaM, we apply SaaM to another text-based method StAR (Wang et al., 2021a), which adopted Siamese network architecture. Table 9 shows the results where StAR+SaaM stands for the StAR model architecture trained by our training framework SaaM. Performance improvement are observed in all evaluation metrics, with significant gain in MRR and Hits@1. There results empirically validate that the model-agnostic SaaM framework can be successfully applied to different text-based methods.
Figure 7: The ratio of triples categorized by relation type for the WN18RR, NELL-995, and FB15k-237 datasets, along with the corresponding Hits@1 performance results for each dataset.
Appendix G Performance of Hybrid In-Batch Negatives
WN18RR
Method MRR Hits@1 Hits@3 Hits@10
Mixed 0.665 0.586 0.722 0.814
SaaM 0.689 0.621 0.731 0.823
FB15k-237
Method MRR Hits@1 Hits@3 Hits@10
Mixed 0.346 0.249 0.388 0.534
SaaM 0.368 0.276 0.401 0.548
Table 10: Performance comparison between Mixed and SaaM, where Mixed replaces half of in-batch triples generated by BRWR with randomly selected triples.
To evaluate the efficacy of constructing a batch exclusively with triples in a subgraph (i.e., SaaM), we compare two sampling methods SaaM and Mixed. Mixed represents a variant that replaces a half of triples in each mini-batch produced by SaaM with randomly selected triples. Table 10 illustrates the performance of SaaM and Mixed. Including random samples deteriorates performances, e.g., a significant drop in Hits@1, which indicates that a mini-batch consisting only of triples within the subgraph is more beneficial.
Triples $\psi_{h}L_{(h,r,t)}$ $L_{(h,r,t)}$
total triples 0.2268 0.1571
false positives 0.4987 0.2834
Table 11: Comparison of average loss values on WN18RR for total triples and false positives in the batch.
Approach WN18RR NELL-995 FB15k-237
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
Embedding-based approach
SANS Ahrabian et al. (2020) 0.216 0.027 0.322 0.509 0.359 0.309 0.381 0.493 0.298 0.203 0.331 0.486
TransE Bordes et al. (2013) 0.239 0.421 0.450 0.510 0.383 0.321 0.400 0.492 0.280 0.193 0.372 0.439
DistMult Yang et al. (2015) 0.435 0.410 0.450 0.510 0.409 0.343 0.427 0.515 0.280 0.195 0.297 0.441
TuckER Balazevic et al. (2019) 0.466 0.432 0.478 0.518 0.423 0.363 0.455 0.536 0.361 0.265 0.391 0.538
RotatE Sun et al. (2019) 0.471 0.421 0.490 0.568 0.411 0.350 0.439 0.522 0.335 0.243 0.374 0.529
KGTuner Zhang et al. (2022) 0.481 0.438 0.499 0.556 0.428 0.371 0.458 0.544 0.345 0.252 0.381 0.534
CompoundE Ge et al. (2023) 0.492 0.452 0.510 0.570 0.434 0.381 0.466 0.550 0.350 0.262 0.390 0.547
CSProm-KG Chen et al. (2023) 0.569 0.520 0.590 0.675 0.422 0.381 0.474 0.535 0.355 0.261 0.389 0.531
Text-based approach
KG-BERT Yao et al. (2019) 0.216 0.040 0.298 0.516 0.199 0.021 0.267 0.488 0.158 0.019 0.232 0.420
MTL-KGC Kim et al. (2020) 0.331 0.203 0.383 0.597 0.288 0.176 0.229 0.411 0.267 0.172 0.298 0.458
StAR Wang et al. (2021a) 0.398 0.238 0.487 0.698 0.419 0.238 0.395 0.433 0.288 0.195 0.313 0.480
HaSa Zhang et al. (2023a) 0.538 0.444 0.588 0.713 0.411 0.310 0.431 0.524 0.304 0.220 0.325 0.483
KG-S2S Chen et al. (2022) 0.572 0.529 0.595 0.663 0.416 0.311 0.421 0.519 0.337 0.255 0.374 0.496
SimKGC Wang et al. (2022a) 0.671 0.580 0.729 0.811 0.425 0.353 0.447 0.531 0.340 0.252 0.365 0.515
GHN Qiao et al. (2023) 0.678 0.596 0.719 0.821 - - - - 0.339 0.251 0.364 0.518
Ensemble approach
StAR(Ensemble) Wang et al. (2021a) 0.520 0.456 0.509 0.707 0.415 0.311 0.402 0.510 0.332 0.229 0.387 0.526
SATKGC-SPW-DW 0.672 0.605 0.714 0.811 0.411 0.310 0.431 0.524 0.304 0.220 0.325 0.483
SATKGC-DW 0.679 0.609 0.723 0.813 0.429 0.361 0.466 0.551 0.351 0.261 0.372 0.525
SATKGC-SPW 0.681 0.612 0.729 0.820 0.431 0.369 0.472 0.556 0.359 0.271 0.393 0.538
SATKGC 0.689 0.621 0.731 0.823 0.433 0.389 0.476 0.560 0.368 0.276 0.401 0.548
Table 12: KGC results for the WN18RR, NELL-995, and FB15k-237 datasets. “SPW” and “DW” refer to shortest path weight and degree weight respectively. The best and second-best performances are denoted in bold and underlined respectively. $\dagger$: numbers are from Qiao et al. (2023).
Appendix H False Positive Analysis
We aim to demonstrate that our contrastive learning with two structure-aware factors selectively penalizes hard negative triples. Table 11 compares $\psi_{h}L_{(h,r,t)}$ and $L_{(h,r,t)}$ on average only for false positives and those for all training triples, where $L_{(h,r,t)}$ in Equation (1) represents loss with only the distance factor $\omega_{h{t_{i}}}$ applied, and $\psi_{h}L_{(h,r,t)}$ in Equation (2) additionally applies the degree factor $\psi_{h}$. The average loss values for the false positives are higher than those for all the triples, which indicates that the structure-aware contrastive learning method severely punishes incorrect predictions for hard negative triples.
Appendix I Correlation of Relation Types and Performances
We investigate the distribution of triples based on their relation types on WN18RR, NELL-995, and FB15k-237. Figure 7 shows that the ratio of triples with the N-N relation type increases in the order of WN18RR, NELL-995, and FB15k-237, while the ratio of triples with the N-1 and 1-N relation types decreases in the same order. In Table 2, both embedding-based and text-based approaches achieve the best results on WN18RR among the three datasets, whereas the performance is worse on FB15k-237. We observe that a high proportion of the N-N relation type and a low proportion of the N-1 and 1-N relation types negatively impact the model performance.
This performance difference between the datasets is larger for a text-based approach than for an embedding-based approach. We speculate that this is because the embedding-based approach randomly initializes the entity and relation embeddings, while the text-based approach uses contextualized text embeddings obtained from PLMs. For the N-N relations where multiple tails can be the correct answer for the same $(h,r)$ pair, the embeddings of these correct tails should be similar. However, PLMs take only text as input, being oblivious of their high similarity. Therefore, these tail embeddings generated by the PLMs might be far apart from each other, so the $(h,r)$ embedding is likely to remain in the middle of these tail embeddings during fine-tuning.
Appendix J Hyperparameter Sensitivity
Figure 8: (a) Impact of varying restart probabilities on the model performance. (b) Impact of varying the number of duplicate entities in a mini-batch on the model performance.
We investigate how restart probability $p_{r}$ in BRWR affects model performance. The hyperparameter $p_{r}$ is associated with the length of the random walk path from the start entity, which in turn influences the occurrence of duplicate entities in a mini-batch. A longer path leads to fewer duplicate entities in the mini-batch. Figure 8(a) illustrates that a lower $p_{r}$ value, encouraging a longer random walk path, leads to higher Hits@1 for WN18RR and FB15k-237. We analyze the impact of duplicate entities in a mini-batch on the model performance. In Figure 8(b), more duplicate entities resulting from higher $p_{r}$ negatively impact on the performance, which highlights the importance of reducing the duplicates in a mini-batch to avoid the performance degradation.
Appendix K Entire Main Results
Table 12 presents the results of all baselines compared with ours.